# inkscapepath=inkscape, inkscapearea=page, clean=true, tex=false

### v1

@@ 1 Introduction
The evolution of open-source text-to-image (TTI) technology has shifted from being a niche area for early adopters to becoming a vast ecosystem driven by increasing investment and commercialization. Estimates by Everypixel (2023) suggest that open-source TTI systems have long outpaced closed-source, API-gated systems in terms of the number of images produced. The technology continues to advance rapidly, with recent innovations allowing users to personalize models through model adapters, tailoring them for specific tasks and creative pursuits such as mimicking a specific art style or reproducing the physiognomy of a particular individual. While model personalizations open up unprecedented opportunities for artists and researchers, they also raise significant ethical and legal concerns, particularly the potential for misuse such as non-consensual deepfakes and disinformation. Platforms like CivitAI [^1] CivitAI https://civitai.com/, (Accessed: 11 November 2024), a pioneering and established hub for developing and sharing open-source TTI models and model adapters, have played a central role in the expansion of AI-generated content and specifically TTI models.
@@
What sets model-sharing hubs apart from other platforms is their function: the content shared on these hubs is not an end product but a tool for generating limitless derivative content. Personalized models, trained on specific concepts, characters, or the physical appearance of a certain individual, can act as powerful multipliers: they enable users to create vast amounts of content that mimics the characteristics of the original training data, which can then be widely shared across social media and other online platforms. This dynamic harbors a high potential for virulence, as its impact is expected to extend far beyond the confines of the communities creating these models. We are particularly concerned about misuse, such as technology-facilitated gender-based violence (TFGBV) (UN Women, 2024), which is facilitated by systems capable of creating entirely fabricated, photorealistic, and believable imagery within seconds. We aim to highlight how model-sharing platforms may exacerbate these concerning developments in their role of fostering environments that are tolerant of abusive behavior, while becoming increasingly hostile towards certain user groups. This trajectory, if left unchecked, may prove more detrimental to open-source innovation than any external regulation. Drawing on a definition of misogyny as a set of hostile mechanisms designed to reinforce patriarchal power structures (Manne, 2017), our objective is to investigate how these dynamics manifest themselves in open-source TTI communities. Specifically, we intend to explore how over-sexualization, objectification, and the erosion of consent in TTI-generated content leads to the proliferation of harmful stereotypes and the marginalization of certain demographic groups, specifically women [^2] We use the terms women and girls to refer to female-presenting individuals. We recognize that gender presentation and identity are diverse, but use simplified terms here for clarity..
@@
In practice, our study seeks to examine the underlying mechanisms contributing to the escalation of NSFW content and the perpetuation of misogynistic tropes within open-source TTI communities, as exemplified by the platform CivitAI. The prevalence of NSFW content is illustrated in Figure 1, which shows a random sample of images taken from CivitAI in March 2024, as categorized by the platform according to different levels of NSFW.
@@
In our study, we perform a detailed analysis of user-generated content, model adapters, and textual training data, in order to uncover how community-driven, tool-driven, and model-infused biases may have intensified within the platform’s ecosystem over the last two years. Our results suggest a pattern of systematic misogyny that contributes to a self-reinforcing cycle of platform degradation, facilitating the spread and normalization of harmful, exclusionary, and exploitative content. Besides acknowledging these issues, we identify potential areas for intervention, such as improving tools used for training models and model adapters, vetting model metadata, and developing and enforcing ethical frameworks. Through these contributions, our work aims to contribute to a more inclusive open-source TTI ecosystem.
2 From Avocado Chairs to Deepfake Porn: The Evolution of Text-to-Image Technologies
In early 2021, OpenAI demonstrated the capabilities of its novel TTI system, DALL·E  and showcased its capabilities with the now legendary depictions of avocado chairs  (OpenAI, 2021), exemplifying the model’s ability to combine distinct concepts to produce novel outputs. OpenAI did not release DALL·E in full, yet made one of its components, the CLIP model (Contrastive Language-Image Pre-Training)  (Radford et al., 2021) publicly available. CLIP was trained on a vast dataset of image–text pairs gathered from the internet, enabling it to associate images with descriptive text at a highly meaningful level. CLIP revolutionized multimodal machine learning because it allowed to directly measure the similarity between images and text, which was utilized in generative models such as DALL·E to create images from prompts with high semantic accuracy.
The open-source community sought to replicate DALL·E’s functionality, by employing CLIP for various implementations of generative image synthesis, such as CLIP + VQGAN (Crowson et al., 2022) or CLIP-Guided Diffusion  (Crowson, 2021). The first implementations of community-developed TTI pipelines were often shared via social media and Discord channels among open-source enthusiasts. These early communities and first implementations were essential for the subsequent later development of more advanced generative methods and were characterized by collaborative efforts, with a focus on creative and artistic applications (Oppenlaender, 2022).
In August 2022, the release of the first Stable Diffusion model by the academic research team CompVis (the Computer Vision and Learning Group at LMU Munich) (Rombach et al., 2022), and companies Runway AI and Stability AI, marked a significant milestone in TTI generation. Stable Diffusion can be regarded as the first open-source TTI foundation model, a large, versatile model trained on vast amounts of data and adaptable to many downstream tasks. The model rapidly became popular because of its high-quality image generation and open-source nature, enabling anyone to use and modify the model. This release also indicated a shift from community-based development to investor-funded commercialized activity in open-source TTI.
After the initial fascination triggered by the rapid technological advances of TTI technologies, attention soon shifted towards understanding their more problematic aspects. As the visual quality of generated images improved, it became increasingly important to address the different issues associated with these models such as copyright violations, challenges in attributing authorship or the various dangers of using synthetic imagery. Most commonly, the critiques of these models focus on the social biases and demographic stereotypes that appear and are perpetuated through AI-generated images, mainly because these models were trained on extremely large, uncurated, web-scraped datasets. Initially examined and discussed through sporadic explorations, these issues have recently developed into a more systematically studied research field, mainly focused on the possibilities of quantifying and evaluating gender and racial biases  (Luccioni et al., 2023; Wu et al., 2024; Naik and Nushi, 2023; Zhang et al., 2024b; Wang et al., 2024; Ghosh and Caliskan, 2023) or cultural stereotypes (Kannen et al., 2024; Cetinic, 2022; Zhang et al., 2024a). Although these studies represent a step forward in better understanding and subsequently mitigating unintentional biases in foundation models, recently such challenges have become more complex with new developments in TTI generation technology, particularly through the emergence of model personalization methods.
These methods, implemented via model adapters, enhance the generative potential of foundation models by providing more refined control over aspects such as style, composition, and subject matter. Model adapters function in conjunction with a base foundation model, adapting it to a specific task. This is accomplished without retraining the entire base model from scratch, requiring far fewer computational resources and fewer training data. Model adapter techniques, such as LoRA (Low-Rank Adaptation for TTI) (Hu et al., 2022) focus on efficiently adapting the model by adding a small number of new trainable parameters, rather than updating the entire set of parameters in the foundation model. In addition to LoRA, there are other model personalization methods such as textual inversion (Gal et al., 2023), where a model learns to associate specific concepts with custom text prompts, allowing for more control over generated content. The process of model personalization became further simplified with the introduction of user-friendly web-based platforms such as Stable-Diffusion-WebUI (automatic1111, 2022) or ComfyUI (comfyanonymous, 2023). These tools allow users to generate state-of-the-art images and videos, train or merge models, and create model adapters through a graphical interface, without requiring advanced technical skills.
The personalization of foundation models can be regarded as a step toward democratizing TTI technology, which, due to vast resource requirements, was previously limited to a few well-funded organizations (Gorwa and Veale, 2024). Although these innovations provide new creative opportunities, they also pose a significant risk of harm, particularly through exploitation and abuse. While traditional deepfake methods typically replace part of a real photo, model adapters such as LoRA can be specifically tailored to reproduce the likeness of any individual in images entirely generated from scratch. These tools, which are capable of portraying individuals in diverse settings based on natural language prompts, enable coordinated attacks that can potentially distort, destroy, or falsely construct a person’s entire digital narrative. The widespread availability of tools for creating ultra-realistic deepfakes is likely to increase digital sexual violations involving compromising images, a form of harm that is already rampant and highly gendered, as female individuals are disproportionately affected by all forms of digital sexual violence  (Viola and Voto, 2023).
The pervasive threat posed by fabricated and compromising digital content is further compounded by the increasing accessibility of models and adapters for generating such material. This shift is partly driven by the emergence of model-sharing hubs. Platforms like Hugging Face [^3] Hugging Face https://huggingface.co/, (Accessed: 17 February 2025), Replicate [^4] Replicate https://replicate.com/, (Accessed: 17 February 2025) and CivitAI have facilitated access to state-of-the-art AI tools by providing inference endpoints for the experimentation with models without the need for technical expertise and expensive hardware, as well as providing the infrastructure to exchange community-created machine learning models, including TTI models.
Although model-sharing platforms have nowadays become central to the distribution and democratization of generative AI models, their role in the propagation of deepfakes, and in particular misogynistic content, has not been systematically studied. While it is evident that such content exists and is shared on these platforms, our exploratory analysis of CivitAI, one of the most popular model-sharing hubs, aims to provide more detailed insight into its prevalence.
@@ 2 From Avocado Chairs to Deepfake Porn: The Evolution of Text-to-Image Technologies
3 Understanding AI Model-Sharing Hubs: A Case Study of CivitAI
In November 2022, the A14Z-funded CivitAI platform was launched, three months after the release of the Stable Diffusion model, as a platform for sharing TTI assets. It has since established itself not only as one of the many sites that offer on-site TTI generations as a service, but primarily as the pioneering and largest platform to exchange community-created TTI personalizations such as Stable Diffusion derivatives and model adapters, as well as other resources used in AI TTI workflows, all arranged in a visually attractive grid layout.
More than 40 million images were posted on CivitAI between late November 2022 and December 2024. The model-sharing hub has gained popularity at a rapid pace: in January 2023, the average number of posted daily images was 864.16. By the end of December 2024, this number had climbed to 178,702.35, a 20,582 % increase. As of 19th January 2025, the site hosts 230.000+ models related to generative image synthesis. Among these, the most widely shared asset type was the LoRA adapter, which accounted for the majority of model adapters on the platform and represented a significant portion of the total exchanges on CivitAI.
Despite being one of the most popular platforms for TTI model sharing, research on CivitAI remains limited. However, some existing studies have revealed the significant presence of NSFW (Not Safe for Work) content. Wei et al. (2024) conducted a large-scale study of CivitAI, uncovering abusive uses of generative models, mainly involving NSFW content and deepfakes. Another study focused on prompt creativity (Palmini et al., 2024), showed a stark increase in the ratio of NSFW content to SFW (Safe for Work) content within AI-generated images posted on CivitAI over a two-year period. Investigative journalism has repeatedly exposed the platform’s role in facilitating the creation of sexually exploitative AI-generated content including potentially illegal material (Maiberg, 2024, 2023). In a broader context of AI platformization, Gorwa and Veale (2024) conducted an analysis of CivitAI, focused on their content moderation policies regarding NSFW and deepfake content, as well as artistic style mimicry issues. They revealed the platform’s antagonistic stance toward artists seeking to remove unauthorized AI models trained on their work. Their findings emphasize the need to go beyond traditional content moderation frameworks and examine the governance challenges posed by open-source generative AI ecosystems. Open-source generative models and sharing platforms present new AI governance challenges, requiring robust technical measures and cross-disciplinary collaboration to overcome the limitations of current regulatory frameworks (Cui and Araujo, 2024). However, deploying effective technical measures to combat emerging issues requires a deeper understanding of the overall dynamics and underlying mechanisms of generative model personalization and sharing environments. Understanding open-source model-sharing hubs requires examining the layered structure through which content is produced and shaped, including generative models as well as the surrounding sociotechnical ecosystem of tools, data, interfaces, and cultural practices. Figure 2 provides a schematic overview of the CivitAI platform, illustrating the various layers of content generation and circulation. Through an exploratory analysis of the CivitAI platform, our study aims to move beyond anecdotal observations and provide data-based conclusions about the systemic challenges and opportunities for improvement within the open-source TTI ecosystem.
A brief visual survey of similar websites suggests that the trends observed on CivitAI are not unique but extend to comparable platforms and may be even more pronounced on less moderated sites. Examples of such platforms include TensorArt [^5] TensorArt https://tensor.art/, (Accessed: 24 November 2024), Pixai [^6] PixAI https://pixai.art/, (Accessed: 12 February 2025), SeaArt [^7] SeaArt https://www.seaart.ai, (Accessed: 12 February 2025), LibLib [^8] LibLib https://www.liblib.art, (Accessed: 13 February 2025). As a prominent platform for exchanging open-source TTI models, CivitAI was selected for this study because of its growing user base, indicating that it is no longer “just another” platform, but rather representative of how uncensored open-source TTI models are currently being created and exchanged. This study does not intend to single out one platform as problematic, but instead aims to highlight systemic issues within the open-source TTI field and propose ways to improve the ecosystem to foster innovation while minimizing downstream harm.
In the following sections we present our data acquisition and categorization process, as well as our quantitative findings, followed by an analysis of the internal mechanisms prevalent on the CivitAI model-sharing platform. Our mixed-methods approach integrates quantitative analyses of the different data categories, with qualitative examinations of user narratives, allowing us to comprehensively assess the platform’s community dynamics. By placing special emphasis on misogynistic elements, we not only quantify their prevalence but also contextualize how gender-related biases manifest within and influence broader community behaviors.
@@ 3.1 Data Collection and Categorization
To uncover patterns that illustrate the evolving dynamics of this prominent model-sharing hub, we collected and analyzed data published on CivitAI over a 2-year period. More specifically, we extracted metadata from millions of images and hundred thousands of models published on CivitAI. Our study covers data from late November 2022 to and including December 2024, obtained via the CivitAI REST API (CivitAI, 2023b). The data collected include metadata from 40,630,560 images [^9] Access to the image metadata dataset can be requested at: https://huggingface.co/datasets/pm-paper-datasets/Civ-40M. and 231,252 models [^10] Access to the model metadata dataset, can be requested at: https://huggingface.co/datasets/pm-paper-datasets/Civ-Models., available on the platform as of 19th January. The model-related metadata includes 9,948 standalone models, such as derivatives of Stable Diffusion, categorized as “checkpoints”, which are fully trained or fine-tuned versions of a model saved for reuse or further modification. And it includes metadata of 221,304 model adapters, such as the aforementioned LoRA or Textual Inversion adapters. In our study, we analyze three categories of data: the metadata associated with images, the metadata of the models shared on the platform, and the metadata related to model training processes.
The image metadata can be divided into two types: content organization metadata and generation metadata. Content organization metadata include details like unique identifiers (CivitAI ID, perceptual hash), creator usernames, timestamps, engagement metrics (e.g. likes), and filtering metadata such as NSFW levels. While NSFW generally refers to content that is inappropriate for workplace settings,covering violent, pornographic, or politically controversial material, on CivitAI the term if effectively synonymous with pornographic content, as non-pornographic NSFW material is quantitatively insignificant. CivitAI classifies image content into five NSFW levels, similar to the Motion Picture Association guidelines for movie ratings  (Motion Picture Association, 2023), which includes the following categories: Browsing Level 1 (PG) is safe for work or SFW; Browsing Level 2 (PG-13) includes revealing clothing, sexy attire, violence, and light gore; Browsing Level 4 (R) consists of adult themes and situations, partial nudity, graphic violence, and death; Browsing Level 8 (X) covers graphic nudity, adult objects, and settings; and Browsing Level 16 (XXX) is reserved for overtly sexual or disturbing graphic content (CivitAI, 2023a). To assign the browsing level, CivitAI uses a combination of Amazon Rekognition (a machine learning service for advanced image and video analysis, including content detection for safety and compliance), unspecified open-source NSFW classification resources, and manual review (CivitAI, 2025).
All information related to image genesis is categorized as generation metadata. This includes parameters set prior to image generation, such as “configuration scale”, which controls how closely the model adheres to the given prompt, and “random seed”, which determines the random noise distribution of the initial image, prior to generation. Additionally, generation metadata includes the models and adapters used for image creation. Specifically, if the image was generated on CivitAI, the metadata contains all CivitAI IDs of the models and adapters used when generating the image. For images generated using a locally run instance, the metadata include hashes denoting the model or models used. An example of how image metadata is structured is included in the Figure 10(a) in the Supplementary Material.
We also collected metadata for models and adapters published on the platform. Model-related metadata include a unique CivitAI ID for each model, with different versions represented as sub-objects, each having its own ID. The metadata also capture download counts, time of creation, engagement metrics, and a POI (person of interest) flag, indicating models and adapters designed to mimic real individuals. Additionally, models are also assigned a NSFW level score by CivitAI, which are not fully explained but appear to indicate NSFW probability in the model’s output. The model metadata include information from the model creation pipeline, preserving details such as the base model used. For this analysis, we focused on the IDs and unique hashes assigned by the framework used for model training to facilitate model identification. For an example of the structure of model metadata, refer to Figure 10(b) in the Supplementary Material.
@@
Finally, we collect the metadata used to identify the auto-tagging system responsible for annotating the training data for model adapters. Training model adapters requires additional data in the form of image-caption pairs, and since manual captioning is time-consuming, auto-tagging systems have become a popular solution for automating caption generation. This metadata were extracted from the downloaded model adapter files (in the *.safetensors format, as described by by321 (2022)). A detailed overview of the data structure can be found in our Supplementary Material Figure 11.

### v2

@@ 1 Introduction
The evolution of open-source text-to-image (TTI) technology has shifted from being a niche area for early adopters to becoming a vast ecosystem driven by increasing investment and commercialization. The technology continues to advance rapidly, with recent innovations allowing users to personalize models through model adapters, tailoring them for specific tasks and creative pursuits such as mimicking a specific art style or reproducing the physiognomy of a particular individual. While model personalizations open up unprecedented opportunities for artists and researchers, they also raise significant ethical and legal concerns, particularly the potential for misuse such as non-consensual deepfakes and disinformation. Platforms like CivitAI [^1] CivitAI https://civitai.com/, (Accessed: 11 November 2024), a pioneering and established hub for developing and sharing open-source TTI models, have played a central role in the expansion of AI-generated content and specifically TTI models.
@@
What sets model-sharing hubs apart from other platforms is their function: the content shared on these hubs is not an end product but a tool for generating limitless derivative content. Personalized models, can act as powerful multipliers as they enable users to create vast amounts of content that mimics the characteristics of the original training data, which can then be widely shared across social media and other online platforms. This dynamic harbors a high potential for virulence, as its impact is expected to extend far beyond the confines of the communities creating these models. We are particularly concerned about misuse, such as technology-facilitated gender-based violence (TFGBV) (UN Women, 2024), which is facilitated by systems capable of creating entirely fabricated, photorealistic, and believable imagery within seconds. We aim to highlight how model-sharing platforms may exacerbate these concerning developments in their role of fostering environments that are tolerant of abusive behavior, while becoming increasingly hostile towards certain user groups. This trajectory, if left unchecked, may prove more detrimental to open-source innovation than any external regulation.
Drawing on a definition of misogyny as a set of hostile mechanisms designed to reinforce patriarchal power structures (Manne, 2017), our objective is to investigate how these dynamics manifest themselves in open-source TTI communities. Specifically, we intend to explore how over-sexualization, objectification, and the erosion of consent in TTI-generated content lead to the proliferation of harmful stereotypes and the marginalization of certain demographic groups, specifically women [^2] We use the terms women and girls to refer to female-presenting individuals. We recognize that gender presentation and identity are diverse, but use simplified terms here for clarity..
@@
In practice, our study seeks to examine the underlying mechanisms contributing to the escalation of NSFW content and the perpetuation of misogynistic tropes within open-source TTI communities, as exemplified by the platform CivitAI. The dominance of NSFW content is illustrated in Figure 1, which shows a random sample of images taken from CivitAI in March 2024, as categorized by the platform according to different levels of NSFW.
@@
In our study, we perform a detailed analysis of user-generated content, model adapters, and textual training data in order to uncover how community-driven, tool-driven, and model-infused biases may have intensified within the platform’s ecosystem over the last two years. Our results suggest a pattern of systematic misogyny that contributes to a self-reinforcing cycle of platform degradation, facilitating the spread and normalization of harmful, exclusionary, and exploitative content.
Besides acknowledging these issues, we identify potential areas for intervention, such as improving tools used for training models and model adapters, vetting model metadata, and developing and enforcing ethical frameworks. Through these efforts, our study aims to contribute to a more inclusive open-source TTI ecosystem.
@@
2 Understanding Model Personalization and Model-Sharing Hubs: A Case Study of CivitAI
In early 2021, OpenAI demonstrated the capabilities of its novel TTI system, DALL·E, and showcased its capabilities with the now legendary depictions of avocado chairs (OpenAI, 2021), exemplifying the model’s ability to combine distinct concepts to produce novel outputs. OpenAI did not release DALL·E in full, yet made one of its components, the CLIP model (Contrastive Language-Image Pre-Training) (Radford et al., 2021), publicly available. CLIP was trained on a large dataset of image–text pairs from the internet, allowing it to associate visuals with descriptive language in a highly meaningful way. This breakthrough transformed multimodal machine learning by enabling direct comparison between images and text, a capability later used in generative models like DALL·E to produce semantically accurate images from prompts.
The open-source community sought to replicate DALL·E’s functionality by employing CLIP for various implementations of generative image synthesis, such as CLIP + VQGAN (Crowson et al., 2022) or CLIP-Guided Diffusion (Crowson, 2021). The first implementations of community-developed TTI pipelines were often shared via social media and Discord channels among open-source enthusiasts. These early communities and first implementations were characterized by collaborative efforts, with a focus on creative and artistic applications (Oppenlaender, 2022).
In August 2022, the release of Stable Diffusion by CompVis, Runway AI, and Stability AI (Rombach et al., 2022) marked a significant milestone as the first open-source TTI foundation model. Its popularity due to high-quality image generation capabilities indicated a shift from community-driven innovation toward investor-funded commercial activity. Concurrently, increased scrutiny arose concerning copyright violations, attribution challenges, and social biases resulting from training these models on uncurated, web-scraped datasets. Systematic research efforts emerged to quantify and evaluate gender, racial, and cultural biases in generated images (Luccioni et al., 2023; Wu et al., 2024; Naik and Nushi, 2023; Zhang et al., 2024b; Wang et al., 2024; Ghosh and Caliskan, 2023; Kannen et al., 2024; Cetinic, 2022; Zhang et al., 2024a).
New developments in TTI technology, notably model personalization through adapters such as LoRA (Low-Rank Adaptation) (Hu et al., 2022), originally developed for language models and later adapted for TTI, and textual inversion (Gal et al., 2023), offered refined control over generated content without retraining entire models. Applications with graphical user interfaces, such as Stable-Diffusion-WebUI (automatic1111, 2022) and ComfyUI (comfyanonymous, 2023), lowered the barrier to entry for TTI, allowing users to personalize models easily and without advanced technical skills. However, model personalization introduced new risks, particularly through the tailored generation of deepfake imagery, exacerbating digital sexual violence disproportionately affecting women (Viola and Voto, 2023). Accessibility increased significantly through model-sharing platforms such as Hugging Face, Replicate, and notably CivitAI, intensifying these potential harms.
Despite the prominence of these platforms, research into their implications remains limited. Existing studies identified prevalent NSFW content, abusive generative practices, and sexually exploitative AI-generated imagery on platforms like CivitAI (Wei et al., 2024; Palmini et al., 2024; Maiberg, 2024, 2023). Broader analyses indicate challenges in governance, highlighting platforms’ antagonistic stance toward artists’ rights and inadequacies of traditional moderation frameworks in addressing these issues (Gorwa and Veale, 2024; Cui and Araujo, 2024).
Our exploratory analysis of CivitAI aims to quantify and contextualize these systemic challenges, emphasizing community dynamics and misogynistic elements. Understanding the layered structure through which content is produced and shaped, including generative models and the surrounding sociotechnical ecosystem, offers valuable insight and supports a more holistic perspective. Figure 2 provides a schematic overview of the CivitAI platform, illustrating the layers of content generation and circulation.
@@ 2.1 Data Collection and Categorization
To illustrate the evolving dynamics of this prominent model-sharing hub, we collected and analyzed image metadata, model metadata, and training process metadata from content published on CivitAI over a 2-year period (late November 2022 to December 2024), using the CivitAI REST API (CivitAI, 2023). The dataset covers 40,630,560 images and 231,252 models as of 19th January [^3] Access available upon request: https://huggingface.co/datasets/pm-paper-datasets. The model metadata includes 9,948 standalone models, commonly referred to as “checkpoints”, capable of generating images independently, and 221,304 adapters (e.g., LoRA, Textual Inversion) that require a checkpoint to function. We analyzed three primary data categories: image metadata, model metadata, and training process metadata.
Image metadata included content organization metadata and generation metadata. Content organization metadata covered identifiers (CivitAI ID, perceptual hash), creator usernames, timestamps, engagement metrics (likes), and NSFW categories. CivitAI classifies images into five NSFW levels: Level 1 (PG) (SFW); Level 2 (PG-13) (revealing clothing, violence); Level 4 (R) (adult themes, partial nudity); Level 8 (X) (graphic nudity); and Level 16 (XXX) (explicit sexual or disturbing content) using Amazon Rekognition and manual review (CivitAI, 2025). Additionally, the image metadata encompassed generation settings such as “configuration scale” and “random seed”, as well as models and adapters used (represented by CivitAI IDs or model hashes). For an example structure, see Figure 10(a) in the Supplementary Material.
Model-related metadata included unique CivitAI IDs, version-specific IDs, download counts, timestamps, engagement metrics, a POI (person of interest) flag indicating mimicry of real individuals. Metadata from model creation pipelines, including base model details and framework-specific hashes, facilitated identification. A detailed structure is shown in Figure 10(b) in the Supplementary Material.
@@
Finally, we extracted metadata written during training of a model, which were used to identify the auto-tagging system responsible for annotating the training data for model personalizations. This metadata was extracted from the downloaded model adapter files (in the *.safetensors format, as described by by321 (2022)). A detailed overview of the data structure can be found in our Supplementary Material, Figure S11.